The dimensions of a rectangular box are measured to be 75 cm, 60 cm,and 40 cm, and each measurement is correct to within 0.2 cm .Use differentials to estimate the largest possible error when the volume of the box is calculated from these measurements.


1
Expert's answer
2022-02-07T10:16:26-0500

volume of the box:

"V=lwh"

where "l=75" cm is length,

"w=60" cm is width,

"h=40" cm is height


largest possible error:

"dV=d(lwh)=whdl+ld(wh)=whdl+l(hdw+wdh)"

where

"dl=dw=dh=0.2" cm


then:

"dV=0.2(60\cdot40+75\cdot40+75\cdot60)=1980" cm3